# Inhomogeneous nonlinear Schrödinger equations with competing singular nonlinearities

### v1

@@ 2.2 Pohozaev Identity and nonexistence results
Suppose $N\geq 2$ and let $0<b<2<q<2^{*}_{b}$. Let $u\in E^{q}_{b}$ be a weak solution for (1.4) where $f:\mathbb{R}^{N}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function given by
@@
for some $\eta_{j}\in[0,2)$ and $1<r_{j}\leq 2^{*}_{\eta_{j}}$ ($r_{j}\in(1,\infty)$ if $N=2$). Then $u\in C^{2}(\mathbb{R}^{N}\backslash\{0\})$ and satisfies the Pohozaev type identity
@@
Let $u\in E_{b}^{q}$ be a weak solution of (1.4). We first show that $u\in C^{2}(\mathbb{R}^{N}\setminus\{0\})$. Denote $h(x,u)=f(x,u)-|x|^{-b}{|u|^{q-2}u}.$ Notice that we have $h(x,u)\in L^{r}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$ for every $r\geq 1$. When $N=2$, this follows directly from the embeddings in Lemma 2.2; when $N\geq 3$, it follows from Lemma B.3 in [31]. Standard Calderón–Zygmund estimates then give $u\in W^{2,r}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$ for all $r\geq 1$, and Schauder estimates yield $u\in C^{2}(\mathbb{R}^{N}\setminus\{0\}).$
To prove the Pohozaev identity, we fix $0<r<R$ and set $A_{r,R}=\{x\in\mathbb{R}^{N}:r<|x|<R\}$. Multiplying the equation (1.4) by $x\cdot\nabla u$ and integrating over $A_{r,R}$, an application of the Green’s identity gives
$\frac{N-2}{2}\int_{A_{r,R}}|\nabla u|^{2}dx+\frac{N-b}{q}\int_{A_{r,R}}\frac{|u|^{q}}{|x|^{b}}dx=\sum_{j=1}^{m}c_{j}\,\frac{N-\eta_{j}}{r_{j}}\int_{A_{r,R}}\frac{|u|^{r_{j}}}{|x|^{\eta_{j}}}dx+\beta(R)-\beta(r),$
@@
Since all integrands belong to $L^{1}(\mathbb{R}^{N})$, we may choose sequences $r_{n}\to 0^{+}$ and $R_{n}\to\infty$ such that $\beta(r_{n})\to 0$ and $\beta(R_{n})\to 0$. Letting $n\to\infty$ yields the identity (2.7). ∎
@@
Assume $0<b<2<q<2^{*}_{b}$ and consider $\eta\in[0,2)$. Let $u\in E^{q}_{b}\cap L^{r}_{\eta}(\mathbb{R}^{N})$ be a solution of the problem
@@ 3 Scaled operators on $E^{q}_{b}(\mathbb{R}^{N})$
    $\exists\ell>0$ such that $\|u_{t}\|=O(t^{\,\ell})$ as $t\to\infty$, uniformly in $u$ on bounded sets.
@@
Suppose $N\geq 2$ and let $0<b<2<q<2^{*}_{b}$. Consider $u_{t}$ as defined in (1.2). The map $E^{q}_{b}\times[0,\infty)\to E^{q}_{b}$ given by $(u,t)\mapsto u_{t}$ is a scaling on $E^{q}_{b}$.
@@
It is straightforward to verify that ($H_{1}$)-($H_{3}$) hold. To prove ($H_{4}$) and ($H_{5}$), recall the notation given in (1.7). Then
@@
which implies ($H_{4}$) and ($H_{5}$) are satisfied with $\ell=(q\delta+b)-N$. Continuity follows then as in [23, Lemma 3.1] as a result of the density of $C^{\infty}_{c}(\mathbb{R}^{N})$ in $E^{q}_{b}$. ∎
@@
An odd potential operator $\tilde{h}$ is said to be a scaled operator provided it maps bounded subsets of $E^{q}_{b}$ into bounded subsets of $(E^{q}_{b})^{*}$ and, for some exponent $\ell\in\mathbb{R}$, satisfies
$\tilde{h}(u_{t})v_{t}=t^{\,\ell}\tilde{h}(u)v\quad\text{for all }u,v\in E^{q}_{b}\text{ and }t\geq 0.$ (3.2)
@@
$\tilde{H}(u_{t})=t^{\,\ell}\tilde{H}(u),\qquad u\in E^{q}_{b},\,t\geq 0.$
@@
Here $2<p<q$ and the parameters satisfy $0<b<a<2$, as specified in (1.3). Both $\mathcal{A}$ and $\mathcal{B}$ fulfill the scaling condition (3.2) with the same exponent $\ell$ as defined in (1.7), which is consistent with assumption ($H_{5}$). Furthermore, the continuous embedding
@@
Notice that if $u\in E^{q}_{b}$ satisfies $\mathcal{A}(u)=\lambda\,\mathcal{B}(u)$, that is, if $u$ is a solution of problem (1.1), then, by the scaling properties of $\mathcal{A}$ and $\mathcal{B}$, its rescaled function $u_{t}$ is also a solution for every $t\geq 0$. Associated with the operators $\mathcal{A}$ and $\mathcal{B}$, we introduce the corresponding potentials
$I,\,J:E^{q}_{b}\longrightarrow\mathbb{R},$
defined by
@@
    For each $u\in E^{q}_{b}\backslash\{0\}$ there is an unique $t>0$ such that $I(tu)=1.$
@@
Multiply (3.6) by q2q-2 and (3.5) by 2b2-b and subtract one from another. Then, divide the resulting equation by 2(Nb)q(N2)2(N-b)-q(N-2), which is nonzero, and we arrive at ($H_{12}$). ∎
@@ 4.1 On the Palais-Smale condition
Here we have $G(x,t)=|x|^{-\eta}H(t)$ in the definition of $\Phi$. Let $c\in\mathbb{R}$ and $(u_{n})$ in $E^{q}_{b}$ a $(PS)_{c}$ sequence for $\Phi,$ namely
@@
To show that it contains a strongly convergent subsequence, we first show that it is bounded in $E^{q}_{b}$. In fact, suppose, arguing by contradiction, that $\|u_{n}\|\to\infty$ for a subsequence. We use the notation (3.7) and set
@@
$\Phi_{\lambda}(u_{n})=\int_{\mathbb{R}^{N}}G(x,u_{n})\,dx+c+o(1).$
Therefore, recalling that
@@
$\Phi_{\lambda}(\tilde{u}_{n})=(\tilde{t}_{n})^{-\ell}\int_{\mathbb{R}^{N}}G(x,u_{n})\,dx+o(1),$ (4.2)
@@
$\Phi^{\prime}_{\lambda}(\tilde{u}_{n})\tilde{u}_{n}={(\tilde{t}_{n})^{-\ell}}\int_{\mathbb{R}^{N}}g(x,u_{n})u_{n}dx+o((\tilde{t}_{n})^{\ell/2}).$ (4.3)
@@
$=\lambda\left(\frac{r}{p}-1\right)\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{p}}{|x|^{a}}dx+(\tilde{t}_{n})^{-\ell}\int_{\mathbb{R}^{N}}\left[rG(x,u_{n})-g(x,u_{n})u_{n}\right]dx+o(1)\leq o(1).$
@@
$\tilde{f}(u)v=\lambda\int_{\mathbb{R}^{N}}\frac{|u|^{p-2}uv}{|x|^{a}}dx+\int_{\mathbb{R}^{N}}g(x,u)vdx.$
@@
This means that $G(x,u)=\frac{1}{2^{*}_{\eta_{1}}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}+\frac{1}{r_{2}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}$ on the definition of functional $\Phi$ in (1.6). Here, the pair $(\eta_{1},2^{*}_{\eta_{1}})$ brings a critical superscaled configuration to problem (1.4), whereas $(\eta_{2},r_{2})$ may be treated in either the subscaled or superscaled regime, even critical. See Definition 1.
@@
Assume $N\geq 3$ and ($\mathcal{H}$). For $f$ given by (4.4), suppose that $\lambda\in\mathbb{R}$, $0\leq\eta_{1}<N-q(N-2)/2$ and $\mu\in\mathbb{R}$. Moreover suppose that the admissible pair $(\eta_{2},r_{2})$ and $\mu$ satisfy one of the following assumptions:
@@
[(i)] For $(\eta_{2},r_{2})$ satisfying i) we have $\ell_{2}\leq\ell$. Since the sequences $(\Phi_{\lambda}(\tilde{u}_{n}))$ and $(\|\tilde{u}_{n}\|_{L^{r_{2}}_{\eta_{2}}})$ are bounded and $\tilde{t}_{n}\to\infty$, it follows from (4.5) that $\tilde{u}_{n}\rightarrow 0$ in $L^{2^{*}_{\eta_{1}}}_{\eta_{1}}(\mathbb{R}^{N})$. This implies that the weak limit of $\tilde{u}_{n}$ in $E^{q}_{b}$ is $v\equiv 0$. Since $b<\eta_{2}$ and $1\leq r_{2}\in(2^{*}_{b,q,\eta_{2}},2^{*}_{\eta_{2}})$, the compactness of the embeddings of $E^{q}_{b}$ in $L^{r_{2}}_{\eta_{2}}(\mathbb{R}^{N})$ and $L^{p}_{a}(\mathbb{R}^{N})$ (see Lemma 2.2 and Corollary 2.3) give also
@@
When $N\geq 3$, observe that $\max\{q,2^{*}_{b,q,\eta}\}=2^{*}_{b,q,\eta}$ whenever $\eta<b$, whereas it equals $q$ if $b\leq\eta<N$. Thus, conditions of the form $r>q$ for an admissible pair $(\eta,r)$ are automatically satisfied whenever $0\leq\eta<b$. Moreover, the condition $\eta<N-\frac{q(N-2)}{2}$ is equivalent to $q<2^{*}_{\eta}$. Since we assume $q<2^{*}_{b}$, this inequality also holds automatically for every $\eta<b$.
@@
Assume ($\mathcal{H}$) and $N\geq 3$. Consider $f(x,u)$ as given in (1.9), with $\lambda\in\mathbb{R}$, $\mu\geq 0$, $\eta_{1}$ satisfying (1.10) and $(\eta_{2},r_{2})$ is a superscaled pair.
If $(r_{2},\eta_{2})$ satisfies (1.11) or (1.12), then the functional $\Phi$ defined in (1.6) satisfies $(PS)_{c}$ for all
@@
If $(r_{2},\eta_{2})$ satisfies (1.13), define $\tilde{S}=\tilde{S}(\mu,\eta_{1},\eta_{2})>0$ such that $\mu S_{\eta_{2}}^{-\frac{2^{*}_{\eta_{2}}}{2}}\tilde{S}^{\frac{2-{\eta_{2}}}{N-2}}+S_{\eta_{1}}^{-\frac{2^{*}_{\eta_{1}}}{2}}\tilde{S}^{\frac{2-\eta_{1}}{N-2}}=1.$ Then, the functional $\Phi$ satisfies the $(PS)_{c}$ condition for all
@@
where we have used that $u_{n}\rightarrow u$ both in $L^{p}_{a}(\mathbb{R}^{N})$ and $L^{r_{2}}_{\eta_{2}}(\mathbb{R}^{N})$. We need now to invoke the Pohozaev Identity given in (2.7) (here is where we need to assume $\eta_{1}<4/N$). Multiply the Pohozaev Identity by $\alpha$ and $\Phi^{\prime}(u)u=0$ by $\beta$, where $\alpha$ and $\beta$ will be chosen later. Then, add the results in (4.16). We get
@@
However, this is a requirement for the pair $(\eta_{2},r_{2})$ to be called superscaled, since
@@
Proof of (ii): Now, let us consider $r_{2}=2^{*}_{\eta_{2}}$ and $\eta_{2}$ satisfying (1.10). From (4.14), using (4.10), we get
@@
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{p}}{|x|^{a}}dx+\left(\frac{1}{2^{*}_{\eta_{1}}}-\frac{1}{2^{*}_{\eta_{2}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx.$
@@
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx+\left(\frac{1}{2^{*}_{\eta_{1}}}-\frac{1}{2^{*}_{\eta_{2}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx.$
@@
The $(PS)_{c}$ condition also holds for levels below the threshold in Lemma 4.3 when $(r_{2},\eta_{2})$ is subscaled, $b<\eta_{2}<2$ and $\mu\leq 0$. Indeed, notice that $(PS)_{c}$ sequences are bounded due to Lemma 4.2. To prove that there exists convergent subsequences, recall that in Lemma 4.3 (the case with only one critical pair) the argument required

### v2

@@ 2.2 Pohozaev Identity and nonexistence results
Suppose $N\geq 2$ and let $0\leq b<2<q<2^{*}_{b}$. Let $u\in E^{q}_{b}$ be a weak solution for (1.4) where $f:\mathbb{R}^{N}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function given by
@@
for some $\eta_{j}\in[0,2)$ and $1<r_{j}\leq 2^{*}_{\eta_{j}}$ ($r_{j}\in(1,\infty)$ if $N=2$). Then $u\in C^{2}(\mathbb{R}^{N}\backslash\{0\})$ and, if $f(\cdot,u)u\in L^{1}(\mathbb{R}^{N})$, it satisfies the Pohozaev type identity
@@
Let $u\in E_{b}^{q}$ be a weak solution of (1.4). We first show that $u\in C^{2}(\mathbb{R}^{N}\setminus\{0\})$. Considering $h(x)=f(x,u(x))-|x|^{-b}{|u(x)|^{q-2}u(x)},$ we verify that $h\in L^{r}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$ for every $r\geq 1$. When $N=2$, this follows directly from the embeddings in Lemma 2.2; when $N\geq 3$, it follows from Lemma B.3 in [31]. Standard Calderón–Zygmund estimates then give $u\in W^{2,r}_{\mathrm{loc}}(\mathbb{R}^{N}\setminus\{0\})$ for all $r\geq 1$, and Schauder estimates yield $u\in C^{2}(\mathbb{R}^{N}\setminus\{0\}).$
To prove the Pohozaev identity, we fix $0<\rho<R$ and set $A_{\rho,R}=\{x\in\mathbb{R}^{N}:\rho<|x|<R\}$. Multiplying the equation (1.4) by $x\cdot\nabla u$ and integrating over $A_{\rho,R}$, an application of the Green’s identity gives
$\frac{N-2}{2}\int_{A_{\rho,R}}|\nabla u|^{2}dx+\frac{N-b}{q}\int_{A_{\rho,R}}\frac{|u|^{q}}{|x|^{b}}dx=\sum_{j=1}^{m}c_{j}\,\frac{N-\eta_{j}}{r_{j}}\int_{A_{\rho,R}}\frac{|u|^{r_{j}}}{|x|^{\eta_{j}}}dx+\beta(R)-\beta(\rho),$
@@
Since all integrands belong to $L^{1}(\mathbb{R}^{N})$, we may choose sequences $\rho_{n}\to 0^{+}$ and $R_{n}\to\infty$ such that $\beta(\rho_{n})\to 0$ and $\beta(R_{n})\to 0$. Letting $n\to\infty$ yields the identity (2.7). ∎
@@
Assume $0\leq b<2<q<2^{*}_{b}$ and consider $\eta\in[0,2)$. Let $u\in E^{q}_{b}\cap L^{r}_{\eta}(\mathbb{R}^{N})$ be a weak solution of the problem
@@ 3 Scaled operators on $E^{q}_{b}(\mathbb{R}^{N})$
    $\exists\sigma>0$ such that $\|u_{t}\|=O(t^{\,\sigma})$ as $t\to\infty$, uniformly in $u$ on bounded sets.
@@
Suppose $N\geq 2$ and let $0\leq b<2<q<2^{*}_{b}$. Consider $u_{t}$ as defined in (1.2). The map $E^{q}_{b}\times[0,\infty)\to E^{q}_{b}$ given by $(u,t)\mapsto u_{t}$ is a scaling on $E^{q}_{b}$.
@@
It is straightforward to verify that ($H_{1}$)-($H_{3}$) hold. To prove ($H_{4}$) and ($H_{5}$), recall the definition of $\ell=(q\delta+b)-N$, given in (1.7). We have
@@
which implies that ($H_{4}$) and ($H_{5}$) are satisfied with $\sigma=\ell$. Continuity follows then as in [23, Lemma 3.1] as a result of the density of $C^{\infty}_{c}(\mathbb{R}^{N})$ in $E^{q}_{b}$. ∎
@@
An odd potential operator $\tilde{h}$ is said to be a scaled operator provided it maps bounded subsets of $E^{q}_{b}$ into bounded subsets of $(E^{q}_{b})^{*}$ and, for $\sigma$ as in ($H_{5}$), it satisfies
$\tilde{h}(u_{t})v_{t}=t^{\,\sigma}\tilde{h}(u)v\quad\text{for all }u,v\in E^{q}_{b}\text{ and }t\geq 0.$ (3.2)
@@
$\tilde{H}(u_{t})=t^{\,\sigma}\tilde{H}(u),\qquad u\in E^{q}_{b},\,t\geq 0.$
@@
Here $2<p<q$ and the parameters satisfy $0\leq b<a<2$, as specified in (1.3). Both $\mathcal{A}$ and $\mathcal{B}$ fulfill the scaling condition (3.2) with $\sigma=\ell$, which is consistent with assumption ($H_{5}$). Furthermore, the continuous embedding
@@
Notice that if $u\in E^{q}_{b}$ satisfies $\mathcal{A}(u)=\lambda\,\mathcal{B}(u)$, that is, if $u$ is a solution of problem (1.1), then, by the scaling properties of $\mathcal{A}$ and $\mathcal{B}$, its rescaled function $u_{t}$ is also a solution for every $t\geq 0$. Associated with the operators $\mathcal{A}$ and $\mathcal{B}$, we introduce the corresponding potentials $I,\,J:E^{q}_{b}\rightarrow\mathbb{R},$ defined by
@@
    For each $u\in E^{q}_{b}\backslash\{0\}$ there is a unique $t>0$ such that $I(tu)=1.$
@@
Multiply (3.6) by q2q-2 and (3.5) by 2b2-b and subtract one from another. Then, divide the resulting equation by 2(Nb)q(N2)2(N-b)-q(N-2), which is nonzero, and we reach ($H_{12}$). ∎
@@ 4.1 On the Palais-Smale condition
Here we have $F(x,t)=\lambda|x|^{-a}|t|^{p}/p+|x|^{-\eta}H(t)$ in the definition of $\Phi$. Let $c\in\mathbb{R}$ and $(u_{n})$ in $E^{q}_{b}$ a $(PS)_{c}$ sequence for $\Phi,$ namely
@@
To show that $(u_{n})$ contains a strongly convergent subsequence, we first show that it is bounded in $E^{q}_{b}$. Arguing by contradiction, suppose that $\|u_{n}\|\to\infty$ for a subsequence. We use the notation (3.7) and set
@@
$\Phi_{\lambda}(u_{n})=\int_{\mathbb{R}^{N}}\frac{H(u_{n})}{|x|^{\eta}}\,dx+c+o(1),$
for $\Phi_{\lambda}$ as defined in (1.6). Therefore, recalling that
@@
$\Phi_{\lambda}(\tilde{u}_{n})=(\tilde{t}_{n})^{-\ell}\int_{\mathbb{R}^{N}}\frac{H(u_{n})}{|x|^{\eta}}\,dx+o(1),$ (4.2)
@@
$\Phi^{\prime}_{\lambda}(\tilde{u}_{n})\tilde{u}_{n}={(\tilde{t}_{n})^{-\ell}}\int_{\mathbb{R}^{N}}\frac{h(u_{n})u_{n}}{|x|^{\eta}}dx+o((\tilde{t}_{n})^{\ell/2}).$ (4.3)
@@
$=\lambda\left(\frac{r}{p}-1\right)\int_{\mathbb{R}^{N}}\frac{|\tilde{u}_{n}|^{p}}{|x|^{a}}dx+(\tilde{t}_{n})^{-\ell}\int_{\mathbb{R}^{N}}\frac{\left[rH(u_{n})-h(u_{n})u_{n}\right]}{|x|^{\eta}}dx+o(1)\leq o(1).$
@@
$\tilde{f}(u)v=\lambda\int_{\mathbb{R}^{N}}\frac{|u|^{p-2}uv}{|x|^{a}}dx+\int_{\mathbb{R}^{N}}\frac{h(u)v}{|x|^{\eta}}dx.$
@@
This means that $F(x,u)=\frac{\lambda}{p}\frac{|u|^{p}}{|x|^{a}}+\frac{1}{2^{*}_{\eta_{1}}}\frac{|u|^{2^{*}_{\eta_{1}}}}{|x|^{\eta_{1}}}+\frac{1}{r_{2}}\frac{|u|^{r_{2}}}{|x|^{\eta_{2}}}$ on the definition of functional $\Phi$ in (1.6). Here, the pair $(\eta_{1},2^{*}_{\eta_{1}})$ brings a critical superscaled configuration to problem (1.4), whereas $(\eta_{2},r_{2})$ may be treated in either the subscaled or superscaled regime, even critical. See Definition 1.
@@
Assume $N\geq 3$ and ($\mathcal{H}$). For $f$ given by (4.4), suppose that $\lambda\in\mathbb{R}$, $\eta_{1}$ satisfies (1.10) and $\mu\in\mathbb{R}$. Moreover suppose that the admissible pair $(\eta_{2},r_{2})$ and $\mu$ satisfy one of the following assumptions:
@@
[(i)] For $(\eta_{2},r_{2})$ satisfying i) we have $\ell_{2}\leq\ell$. Since the sequences $(\Phi_{\lambda}(\tilde{u}_{n}))$ and $(\|\tilde{u}_{n}\|_{L^{r_{2}}_{\eta_{2}}})$ are bounded and $\tilde{t}_{n}\to\infty$, it follows from (4.5) that $\tilde{u}_{n}\rightarrow 0$ in $L^{2^{*}_{\eta_{1}}}_{\eta_{1}}(\mathbb{R}^{N})$. This implies that the weak limit of $\tilde{u}_{n}$ in $E^{q}_{b}$ is $v\equiv 0$. Since $b<\eta_{2}$ and $1<r_{2}\in(2^{*}_{b,q,\eta_{2}},2^{*}_{\eta_{2}})$, the compactness of the embeddings of $E^{q}_{b}$ in $L^{r_{2}}_{\eta_{2}}(\mathbb{R}^{N})$ and $L^{p}_{a}(\mathbb{R}^{N})$ (see Lemma 2.2 and Corollary 2.3) give also
@@
When $N\geq 3$, observe that $\max\{q,2^{*}_{b,q,\eta}\}=2^{*}_{b,q,\eta}$ whenever $\eta<b$, whereas it equals $q$ if $b\leq\eta<N$. Thus, conditions of the form $r>q$ for an admissible pair $(\eta,r)$ are automatically satisfied whenever $0\leq\eta<b$. Moreover, the condition $\eta<N-{q(N-2)}/{2}$ is equivalent to $q<2^{*}_{\eta}$. Since we assume $q<2^{*}_{b}$, this inequality also holds automatically for every $\eta<b$.
@@
Assume ($\mathcal{H}$) and $N\geq 3$. Consider $f(x,u)$ as given in (1.9), with $\lambda\in\mathbb{R}$, $\mu\geq 0$, $\eta_{1}$ satisfying (1.10) and assume that $(\eta_{2},r_{2})$ is a superscaled pair.
If $(\eta_{2},r_{2})$ satisfies (1.11) or (1.12), then the functional $\Phi$ defined in (1.6) satisfies $(PS)_{c}$ for all
@@
If $(\eta_{2},r_{2})$ satisfies (1.13), define $\tilde{S}=\tilde{S}(\mu,\eta_{1},\eta_{2})>0$ such that $\mu S_{\eta_{2}}^{-\frac{2^{*}_{\eta_{2}}}{2}}\tilde{S}^{\frac{2-{\eta_{2}}}{N-2}}+S_{\eta_{1}}^{-\frac{2^{*}_{\eta_{1}}}{2}}\tilde{S}^{\frac{2-\eta_{1}}{N-2}}=1.$ Then, the functional $\Phi$ satisfies the $(PS)_{c}$ condition for all
@@
where we have used that $u_{n}\rightarrow u$ both in $L^{p}_{a}(\mathbb{R}^{N})$ and $L^{r_{2}}_{\eta_{2}}(\mathbb{R}^{N})$. We need now to invoke the Pohozaev Identity given in (2.7). Multiply the Pohozaev Identity by $\alpha$ and $\Phi^{\prime}(u)u=0$ by $\beta$, where $\alpha$ and $\beta$ will be chosen later. Then, add the results in (4.16). We get
@@
However, this is precisely the requirement for the pair $(\eta_{2},r_{2})$ to be called superscaled, since
@@
Proof of (ii): Now, let us consider $\eta_{2}$ satisfying (1.10) and $r_{2}=2^{*}_{\eta_{2}}$. From (4.14), using (4.10), we get
@@
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{p}}{|x|^{a}}dx+\mu\left(\frac{1}{2^{*}_{\eta_{1}}}-\frac{1}{2^{*}_{\eta_{2}}}\right)\int_{\mathbb{R}^{N}}\frac{|u_{n}|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx.$
@@
$-\lambda\left(\frac{1}{p}-\frac{1}{2^{*}_{\eta_{1}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{p}}{|x|^{a}}dx+\mu\left(\frac{1}{2^{*}_{\eta_{1}}}-\frac{1}{2^{*}_{\eta_{2}}}\right)\int_{\mathbb{R}^{N}}\frac{|u|^{2^{*}_{\eta_{2}}}}{|x|^{\eta_{2}}}dx.$
@@
The $(PS)_{c}$ condition also holds for levels below the threshold in Lemma 4.3 when $(\eta_{2},r_{2})$ is subscaled, $b<\eta_{2}<2$ and $\mu\leq 0$. Indeed, notice that $(PS)_{c}$ sequences are bounded due to Lemma 4.2. To prove that there exist convergent subsequences, we use Corollary 2.4, follow the same steps of the proof above, and recall that, in the case with only one critical pair, the argument required